Summary information from the financial statements of two companies competing in the same industry follows.

Barco
Company
Kyan
Company
Barco
Company
Kyan
Company
Data from the current year-end balance sheets Data from the current year’s income statement
Assets Sales $ 780,000 $ 908,200
Cash $ 21,500 $ 33,000 Cost of goods sold 591,100 634,500
Accounts receivable, net 35,400 60,400 Interest expense 8,100 14,000
Current notes receivable (trade) 9,400 8,000 Income tax expense 14,992 25,073
Merchandise inventory 84,740 128,500 Net income 165,808 234,627
Prepaid expenses 5,400 6,900 Basic earnings per share 4.15 5.43
Plant assets, net 320,000 308,400 Cash dividends per share 3.71 3.95
Total assets $ 476,440 $ 545,200
Beginning-of-year balance sheet data
Liabilities and Equity Accounts receivable, net $ 27,800 $ 53,200
Current liabilities $ 60,340 $ 101,300 Current notes receivable (trade) 0 0
Long-term notes payable 84,800 101,000 Merchandise inventory 63,600 107,400
Common stock, $5 par value 200,000 216,000 Total assets 388,000 412,500
Retained earnings 131,300 126,900 Common stock, $5 par value 200,000 216,000
Total liabilities and equity $ 476,440 $ 545,200 Retained earnings 113,892 62,913

1a. For both companies compute the (a) current ratio, (b) acid-test ratio, (c) accounts (including notes) receivable turnover, (d) inventory turnover, (e) days’ sales in inventory, and (f) days’ sales uncollected. (Do not round intermediate calculations.)
1b. Identify the company you consider to be better short-term credit risk.
Answer #1
Ans. 1 a
*Calculations for total current assets:
Assets: Bacro company Kyan company
Cash $21,500 $33,000
Accounts receivables $35,400 $60,400
Current notes receivable $9,400 $8,000
Merchandise inventory $84,740 $128,500
Prepaid expenses $5,400 $6,900
Total current assets (a) $156,440 $236,800
Ans. 1 a - a Current ratio   =   Total current assets / Total current liabilities
Bacro $156,440 / $60,340 2.59 : 1
Kyan $236,800 / $101,300 2.34 : 1
Ans. 1 a - b Acid test ratio   =   (Total current assets - Inventory - Prepaid expenses) / Total current liabilities
Bacro ($156,440 - $84,740 - $5,400) / $60,340 1.099 : 1
Kyan ($236,800 - $128,500 - $6,900) / $101,300 1.000 : 1
Ans. 1 a - c Accounts receivable turnover = Net credit sales / Average receivables
Bacro $780,000 / $36,300 21.49 times
Kyan $908,200 / $60,800 14.94 times
*Average receivable = (Beginning accounts and notes receivables + Ending accounts and notes receivables) / 2
Bacro ($35,400 + $9,400 + $27,800 + $0) / 2 $36,300
Kyan ($60,400 + $8,000 + $53,200 + $0) / 2 $60,800
Ans. 1 a - d Inventory turnover = Cost of goods sold / Average Inventory
Bacro $591,100 / $74,170 7.97 times
Kyan $634,500 / $117,950 5.38 times
*Average inventory = (Beginning inventory + Ending inventory) / 2
Bacro ($63,600 + $84,740) / 2 $74,170
Kyan ($107,400 + $128,500) / 2 $117,950
Ans. 1 a - e Days sales in inventory = No. of days in year / Inventory turnover ratio
Bacro 365 / 7.97 45.80 days
Kyan 365 / 5.38 67.84 days
Ans. 1 a - f Days sales uncollected = No. of days in year / Accounts receivable turnover ratio
Bacro 365 / 21.49 16.98 days
Kyan 365 / 14.94 24.43 days
Ans. 1 b Both companies have approximately equal current and quick ratios so these ratios can not be used
for the measure of short term credit risk.
But, Bacro company takes less time to collect its receivables and turn its inventory so this company
is considered to have a better short term credit risk.